# Self-Training: A Survey

### v5

@@ 3.2 Consistency-based approaches
Other studies refined the Mean-Teacher approach with a data-augmentation technique by combining two images with random patches to improve prediction consistency [33, 98]. More recently, [26] provide a two-stage method to reduce label propagation errors; where in the first phase, the gradients of the student loss are computed and utilized to update the teacher. In the second stage, the teacher assigns pseudo-labels which are then utilized to train the current student.
@@ 3.3 Self-supervised Learning
Common approaches in self-supervised learning include predicting missing parts of an image [53], predicting the order of shuffled image patches or their orientation [85], reconstructing corrupted images [28], filling in missing words in a sentence [25], or predicting future frames in a video sequence [80]. These pretext tasks encourage the model to capture meaningful representations of the input data, which can then be used for various downstream tasks, such as image classification, object detection, or natural language processing.
@@ 4.1 Natural Language Processing
Other current self-training techniques in NLP are mostly built on the concept of co-training and employ two base classifiers that are learned over each other’s predictions. In this line, [97] proposed a Named Entity Recognition (NER) strategy that consists in automatically detecting and classifying named entities, with a first NER model trained on labeled training data serving as a teacher to predict the probability distribution of entity labels for each token in the unlabeled set. The pseudo-labeled data with such soft labels are then used to train a student NER model for the unlabeled set and the process of pseudo-labeling and training is repeated until convergence as in co-training. For the task of Relation Extraction (RE) which consists in obtaining a predefined semantic relation between two entities in a given sentence, [103] proposed an approach which classifies the pseudo-labeled instances generated from a teacher into confident, ambiguous and hard sets. In the training of the student model, the confident and ambiguous instances are subsequently interpreted as positive and set-negatives observations, respectively.
Lately, [61] proposed an approach to leverage the power of language models that have been pre-trained on large corpora of text to generate pseudo-labels for unlabeled text data. The pseudo-labeled data along with a smaller set of labeled data are then used to train and fine-tune the text classifier, and the process of assigning pseudo-labels and retraining of the classifier is repeated until convergence. A challenge that arises when using a single base classifier in self-training for NLP tasks is to minimize the impact of label noise in the pseudo-labeling policy. To cope with this problem, [105] devised a bootstrapping technique for semantic role labeling that consists randomly selecting a subset of the most confident samples for pseudo-labeling. In the same vein, for the sentiment analysis task, [39] advocated selecting the top most confident samples for pseudo-labeling. However, as we shall see in the next section, the use of a fixed threshold in the pseudo-labeling policy may be suboptimal in general.
@@ 4.2 Computer Vision
[88] proposed a self-training approach called FixMatch that combines consistency regularization with a confidence-based mechanism to select high-confidence pseudo-labeled examples for training. The algorithm applies to the same image two different data augmentations procedures, called weak (flip-and-shift) and strong (more heavy distortions) augmentations. As in the previous case, these perturbations helps to increase diversity and improve the model’s robustness on the unlabeled images. The authors introduce a consistency loss term that encourages the consistency between the model’s hard output of the weakly-augmented version and the model’s soft output of the strongly-augmented version of the same unlabeled image. They demonstrate that the model learns to provide more trustworthy and accurate results by minimizing the discrepancy between these predictions. In order to decrease the influence of possibly inaccurate pseudo-labels on the learning process, the loss is evaluated only on those unlabeled data from the batch that have the confidence higher than a fixed threshold.
@@ 4.3.1 Speech Recognition
Other approaches integrate domain-specific language models in the the pseudo-label generation process in order to ensure that the generated labels adhere to the linguistic nuances and terminologies of the domain. In this line, [47] introduced a self-training approach, with one base classifier combined with a language model for pseudo-labeling. Their approach involves implementing tailored filtering methods designed to address common errors arising from sequence-to-sequence models, alongside an inventive ensemble technique for enhancing the breadth of pseudo-label variations. Building upon this idea, [100] showcased that the synergy between self-training and unsupervised pre-training using wav2vec 2.0 [7] offers mutual benefits across diverse scenarios involving labeled and unlabeled data.
@@ 4.3.2 Anomaly Detection
Leveraging domain knowledge to mitigate label noise in pseudo-labels within self-training approaches has also been considered in anomaly detection. In this case, the understanding of the anomaly patterns and characteristics specific to the domain are incorporated in the model. In this regard, [55] identified common anomaly types, their features and potential sources of noise and [74] performed time domain analysis. Also, [29] created features that capture domain-specific information for video anomaly detection. It was demonstrated that these features highlight the crucial elements for anomaly detection in videos, leading to their utilization for enhancing the pseudo-labeling phase within self-training. Alternate strategies focus on simulating anomalies within the unlabeled dataset using domain knowledge. This aids the model in learning from a broad spectrum of anomalies, mitigating the potential of becoming overly specialized in a particular anomaly type [73].
@@ 4.3.3 Genomics and proteomics
Additionally, [27] employed self-training in conjunction with bioinformatic tools in silico to anticipate secreted and protective proteins. This was done to eliminate pseudo-label errors from the identified P. falciparum SEPs obtained through proteomics experiments and to anticipate new SEPs within the P. falciparum proteome.
@@
The key observations made in these applications reveal that, in pseudo-labeling, employing fixed thresholds often yields suboptimal outcomes, underscoring the importance of dynamic thresholding for optimal results. Furthermore, enhancing pseudo-label noise improves both generalization and class differentiation.
5 Conclusion and Perspectives
@@
First, we discussed the various strategies for selecting unlabeled samples for pseudo-labeling that have been proposed. We emphasized the significance of considering margin distributions across unlabeled data as a pivotal factor in the development of these strategies. Next, we provided an overview of the diverse variants of self-training explored in the literature, along with relevant approaches. Furthermore, we examined recent theoretical advancements in this research domain and outlined the principal characteristics of self-training employed in several widely recognized applications.
@@
Recent research emphasizes the potential of exploring self-training methods from a theoretical standpoint, particularly in addressing the challenge of training a final classifier on data with noisy labels [40]. It has also been demonstrated that accurately estimating the confidence of pseudo-labels is crucial for effective self-training [68]. Therefore, theoretically establishing the correlation between performance and the level of uncertainty in pseudo-labeling could be a valuable direction for future research, especially in analyzing self-training within the context of learning problems affected by domain shifts.

### v6

@@ 3.2 Consistency-based approaches
More recently, [22] provide a two-stage method to reduce label propagation errors; where in the first phase, the gradients of the student loss are computed and utilized to update the teacher. In the second stage, the teacher assigns pseudo-labels which are then utilized to train the current student.
Also, [27] proposed a new contrastive semi-supervised learning approach for data-efficient language modeling. The method uses hard prompts for sentence representation and integrates prompt-based pseudo labeling with a mask language model through a contrast loss, which pulls together similar samples and pushes apart dissimilar ones. Additionally, mask consistency training is employed to align word predictions from weak and strong augmentations. This approach improves model generalization and leverages unlabeled data for few-shot text classification,
@@ 3.3 Self-supervised Learning
Common approaches in self-supervised learning include predicting the order of shuffled image patches or their orientation [91], reconstructing corrupted images [23], filling in missing words in a sentence [8], or predicting future frames in a video sequence [69]. These pretext tasks encourage the model to capture meaningful representations of the input data, which can then be used for various downstream tasks, such as image classification, object detection, or natural language processing.
@@ 4.1 Natural Language Processing
Other current self-training techniques in NLP are mostly built on the concept of co-training and employ two base classifiers that are learned over each other’s predictions. In this line, [89] proposed a Named Entity Recognition (NER) strategy that consists in automatically detecting and classifying named entities, with a first NER model trained on labeled training data serving as a teacher to predict the probability distribution of entity labels for each token in the unlabeled set. The pseudo-labeled data with such soft labels are then used to train a student NER model for the unlabeled set and the process of pseudo-labeling and training is repeated until convergence as in co-training. For the task of Relation Extraction (RE) which consists in obtaining a predefined semantic relation between two entities in a given sentence, [93] proposed an approach which classifies the pseudo-labeled instances generated from a teacher into confident, ambiguous and hard sets. In the training of the student model, the confident and ambiguous instances are subsequently interpreted as positive and set-negatives observations, respectively. Based on these ideas, [50] proposed several self-training methods for NLP, demonstrating that neural language models can be enhanced through synthetic data generation strategies such as pseudo-labeling, prototypical label embeddings, and synthetic text generation. These methods were shown to significantly improve performance in tasks like dialog response retrieval, dialog policy learning, and question answering, achieving state-of-the-art results and reducing dependence on human annotation.
Lately [97] proposed a self-training approach that reduces the annotation cost for structured label spaces using active learning. The method leverages partial annotation by selecting only the most informative substructures for annotation and incorporates the model’s automatic predictions as pseudo-labels for unannotated sub-structures. By employing an error estimator to adaptively decide the partial selection ratio, the approach effectively combines partial annotation with self-training, reducing annotation cost across four structured prediction tasks compared to strong full annotation baselines.
@@ 4.2 Computer Vision
[77] proposed a self-training approach called FixMatch that combines consistency regularization with a confidence-based mechanism to select high-confidence pseudo-labeled examples for training. The algorithm applies to the same image two different data augmentations procedures, called weak (flip-and-shift) and strong (more heavy distortions) augmentations.
As in the previous case, these perturbations helps to increase diversity and improve the model’s robustness on the unlabeled images. The authors introduce a consistency loss term that encourages the consistency between the model’s hard output of the weakly-augmented version and the model’s soft output of the strongly-augmented version of the same unlabeled image. They demonstrate that the model learns to provide more trustworthy and accurate results by minimizing the discrepancy between these predictions. In order to decrease the influence of possibly inaccurate pseudo-labels on the learning process, the loss is evaluated only on those unlabeled data from the batch that have the confidence higher than a fixed threshold.
@@ 4.3.1 Speech Recognition
Other approaches integrate domain-specific language models in the the pseudo-label generation process in order to ensure that the generated labels adhere to the linguistic nuances and terminologies of the domain. In this line, [38] introduced a self-training approach, with one base classifier combined with a language model for pseudo-labeling. Their approach involves implementing tailored filtering methods designed to address common errors arising from sequence-to-sequence models, alongside an inventive ensemble technique for enhancing the breadth of pseudo-label variations. Furthermore [76] presents a self-training approach for automatic speech recognition in low-resource settings, specifically focusing on the Punjabi language. The proposed method generates highly accurate pseudo-labels for unlabeled Punjabi speech, resulting in a significant improvement in word error rate compared to state-of-the-art approaches.
@@ 4.3.2 Anomaly Detection
Leveraging domain knowledge to mitigate label noise in pseudo-labels within self-training approaches has also been considered in anomaly detection. In this case, the understanding of the anomaly patterns and characteristics specific to the domain are incorporated in the model. In this regard, [64] identified common anomaly types, their features and potential sources of noise and performed time domain analysis. Also, [80] proposed a method for video anomaly detection that employs an unsupervised approach using isolation trees and deep features. The method generates initial anomaly and dynamicity scores, which are refined using a cross-branch feed-forward network based on the I3D architecture. It then incorporates a self-training strategy to iteratively improve the model’s performance by pseudolabeling and retraining on the most confident samples, achieving competitive accuracy on popular datasets like UCF-Crime, CCTV-Fights, and UBI-Fights. This approach combines appearance and motion information to enhance the quality of anomaly evidence, making it more practical for real-world applications where annotation is burdensome. Alternate strategies focus on simulating anomalies within the unlabeled dataset using domain knowledge. This aids the model in learning from a broad spectrum of anomalies, mitigating the potential of becoming overly specialized in a particular anomaly type [63]. [47] also proposed a weakly supervised Video Anomaly Detection approach using Multi-Sequence Learning (MSL) and a hinge-based MSL ranking loss to reduce selection errors by optimizing sequences of snippets. The method employs a Transformer-based MSL network to learn video-level and snippet-level anomaly scores, refining them through a self-training strategy that gradually reduces sequence length. This approach achieves significant improvements on datasets like ShanghaiTech, UCF-Crime, and XD-Violence.
@@ 4.3.3 Genomics and proteomics
[45] proposed a self-training subspace clustering algorithm based on adaptive confidence for gene expression data. The proposed algorithm enhances the discriminative property of gene expression data using low-rank representation with distance penalty and introduces a semi-supervised clustering objective function with label confidence. To mitigate the negative impact of mislabeled data, an adaptive adjustment strategy based on the gravitational search algorithm is employed.
@@
The key observations made in these applications reveal that, in pseudo-labeling, employing fixed thresholds often yields suboptimal outcomes, underscoring the importance of dynamic thresholding for optimal results. Furthermore, enhancing pseudo-label noise improves both generalization and class differentiation. In Appendix A, we will show the impact of dynamic thresholding on pseudo-labeling across general benchmarks and examine the noise considerations in two image classification collections studied in [15].
5 Conclusion, Limitations and Perspectives
@@
First, we discussed the various strategies for selecting unlabeled samples for pseudo-labeling that have been proposed. We emphasized the significance of considering margin distributions across unlabeled data as a pivotal factor in the development of these strategies. Next, we provided an overview of the diverse variants of self-training explored in the literature, along with relevant approaches. Furthermore, we examined recent theoretical advancements in this research domain and outlined the principal characteristics of self-training employed in several widely recognized applications. Lastly, we explored the impact of fundamental aspects of self-training on a range of benchmark datasets.
Limitations
Despite the promising results and widespread adoption of self-training, several limitations remain. One significant challenge is the sensitivity of self-training methods to the quality of pseudo-labels. Incorrect or noisy pseudo-labels can propagate errors through the training process, leading to suboptimal performance. Additionally, the effectiveness of self-training can be highly dependent on the initial labeled dataset. If the initial dataset is not representative or is too small, the self-training process may not yield significant improvements.
Another limitation is the computational cost associated with iterative pseudo-labeling and retraining. Self-training methods often require multiple iterations, which can be time-consuming and resource-intensive, especially for large-scale datasets. Furthermore, the selection of appropriate thresholds for pseudo-labeling remains a critical and often challenging task, as it requires a delicate balance between the quantity and quality of pseudo-labels.
Future Work
To address these limitations, future research should focus on developing more robust and efficient self-training algorithms. This could involve exploring adaptive thresholding techniques that dynamically adjust the confidence thresholds based on the model’s learning status. Additionally, incorporating uncertainty quantification methods to estimate the reliability of pseudo-labels could help mitigate the impact of noisy labels.
Another promising direction is the integration of self-training with other semi-supervised learning techniques, such as consistency regularization and multi-view training, to leverage their complementary strengths. Furthermore, extending self-training to new domains, such as medical imaging and industrial time-series data, could open up new avenues for practical applications.
@@
Also, recent research emphasizes the potential of exploring self-training methods from a theoretical standpoint, particularly in addressing the challenge of training a final classifier on data with noisy labels [32]. It has also been demonstrated that accurately estimating the confidence of pseudo-labels is crucial for effective self-training [58]. Therefore, theoretically establishing the correlation between performance and the level of uncertainty in pseudo-labeling could be a valuable direction for future research, especially in analyzing self-training within the context of learning problems affected by domain shifts.
Appendix A Empirical Study
In this section, we will evaluate the effectiveness and performance of the self-training algorithm. This assessment will focus on two key features discussed in the preceding sections: the handling of noise in the pseudo-labeling phase and the impact of threshold selection. Our primary focus will be on scenarios with sufficient labeled training data, allowing for the development of an initial supervised complex model, and scenarios with severely limited labeled training data, where using complex baseline classifiers like deep learning models is not feasible.
A.1 Noise Account
We first consider the case where the initial labeled training set allows to train deep neural networks and examine the effects of taking into account noise in the pseudo-labeling process along with the dynamic selection of the threshold on CIFAR-10 and CIFAR-100 [86]. Both datasets contain 32x32 pixel RGB images belonging to respectively 10 and 100 classes; 50000 examples are used for training and 10000 samples for test.
We consider the debiased self-training approach (DST) [15] to address the presence of noise in pseudo-labeling, in conjunction with the FlexMatch method [95] for the dynamic threshold determination in pseudo-labeling. As outlined in Section 2.3, DST involves training a dedicated head on pseudo-labeled examples, allowing the model to implicitly capture and account for noise inherent in the pseudo-labels.
For FlexMatch, we followed the same experimental protocol than [95]. In this case, Wide ResNet (WRN) was used as the base classifier in self-training. Parameter learning was accomplished using stochastic gradient descent (SGD) with a momentum coefficient of 0.9. The initial rate was set to $\eta_{0}=0.03$ with a cosine learning rate decay schedule as $\eta=\eta_{0}\cos(7\pi t/16T)$, where $t$ denotes the current training step and $T$ is the total training step set at $2^{20}$. Additionally, exponential moving averaging with a momentum of 0.999 was implemented and the batch size for labeled data was fixed to 64. For DST, we used the code made available by the authors [^1] https://github.com/thuml/Debiased-Self-Training.
We compared FlexMatch with and without the DST approach denoted respectively by FM and FM+DST. We also compared self-training with WRN trained in fully supervised manner. Each experiment was repeated 5 times by changing the seed at each time. Figure 2 presents the average accuracy of different models on the test set for the same number of initial labeled training samples per class within the set $\{4,10,20,50\}$ for both datasets. In both datasets, considering label noise within pseudo-labels leads to improved performance, with the improvement being more pronounced in the case of CIFAR-100.
404040401001001001002002002002005005005005009090909092929292949494949696969698989898100100100100# of labeled training data, $m$Accuracy (%)CIFAR-10SupervisedFM+DSTFM 4004004004001,00010001{,}0001 , $0002{,}$ $0005{,}$ 00050505050606060607070707080808080# of labeled training data, $m$Accuracy (%)CIFAR-100SupervisedFM+DSTFM
Figure 2: Comparisons in terms of Accuracy on CIFAR-10 and CIFAR-100 for a varying number of labeled training data. “Supervised” refers to the fully supervised learning ($m=50000$, $u=0$).
In CIFAR-100, classes are structured into 20 superclasses, each comprising 5 related classes, addressing noise in this more complex task aids in class differentiation and enhances the model’s ability to generalize. It is worth noting that with a greater number of initial labeled training examples, the gap between the FM and FM+DST approaches narrows, as the model becomes more proficient with the increased labeled data and makes fewer errors in pseudo-labeling.
A.2 The impact of threshold selection
We now study the effect of selecting automatically the threshold for pseudo-labeling on 9 publicly available data sets proposed for semi-supervised learning [^2] https://archive.ics.uci.edu/ml/index.php. The characteristics of these datasets are presented in Table 2. It is worth noting that certain datasets contain only a limited number of labeled training examples, comprising just a few hundred instances and accounting for less than 1% of the total training examples. This condition underscores the suitability of employing complex base classifiers.
Data set # of labeled examples # of unlabeled examples Dimension # of classes
$m$ $u$ $d$ $K$
Vowel 99 891 10 11
Protein 129 951 77 8
PageBlocks 1094 4379 10 5
Isolet 389 7408 617 26
HAR 102 10197 561 6
Pendigits 109 10883 16 10
Letter 400 19600 16 26
Fashion 175 69825 784 10
MNIST 175 69825 784 10
Table 2: Characteristics of data sets used in our experiments, $d$ and $K$ correspond to respectively the dimension of the input space and the number of classes.
In the experimentation, Random Forest was employed instead using the scikit-learn implementation [61] with 200 trees of maximum depth while leaving other parameters at their default values. The primary objective was to assess and contrast the classifier’s performance in two scenarios: the supervised scenario (denoted by RF) and the self-training scenario where pseudo-labeling is automatically conducted following the approach introduced by [24] [^3] https://github.com/vfeofanov/trans-bounds-maj-vote (denoted by PL). Additionally, we investigated the impact of setting the pseudo-labeling threshold at predefined values from the set $\theta\in\{0.5,0.7,0.9\}$ (denoted by PLθ).
The automatic pseudo-labeling strategy selects the threshold which minimizes the bound of the error of the Random Forest classifier over the unlabeled training samples.
Results are resumed in Table 3. Experiments are repeated 20 times by choosing randomly the labeled training examples, and indicates that performance is statistically worse than the best result, shown in bold, according to the Wilcoxon rank-sum test.
[…]
Vowel .586 $\pm$ .028 .489±limit-fromsuperscript.$489^{\downarrow}\pm.$016 ${}^{\downarrow}\pm$034 ${}^{\downarrow}\pm$028 .586 $\pm$ .026
Protein ${}^{\downarrow}\pm$032 ${}^{\downarrow}\pm$024 .687 $\pm$ .036 .724 $\pm$ .018 .781 $\pm$ .034
PageBlocks .965 $\pm$ .003 ${}^{\downarrow}\pm$003 .964 $\pm$ .004 .965 $\pm$ .002 .966 $\pm$ .002
Isolet .$016.854^{\downarrow}\pm.$016 .$018.648^{\downarrow}\pm.$018 .7 $\pm$ .04 .861 $\pm$ .08 .875 $\pm$ .014
HAR .851 $\pm$.024 ${}^{\downarrow}\pm$04 ${}^{\downarrow}\pm$041 ${}^{\downarrow}\pm$035 .854 $\pm$ .026
Pendigits ${}^{\downarrow}\pm$022 ${}^{\downarrow}\pm$022 ${}^{\downarrow}\pm$036 ${}^{\downarrow}\pm$024 .884 $\pm$ .022
Letter .711 $\pm$ .011 ${}^{\downarrow}\pm$011 ${}^{\downarrow}\pm$015 .673 ${}^{\downarrow}\pm$015 .717 $\pm$ .013
Fashion .718 $\pm$ .022 ${}^{\downarrow}\pm$014 ${}^{\downarrow}\pm$04 ${}^{\downarrow}\pm$014 .723 $\pm$ .023
MNIST ${}^{\downarrow}\pm$015 ${}^{\downarrow}\pm$012 ${}^{\downarrow}\pm$055 ${}^{\downarrow}\pm$045 .857 $\pm$ .013
Table 3: Classification performance using the accuracy score on 9 publicly available data set. Best results are shown in bold and the sign shows if the performance is statistically worse than the best result on the level 0.01 of significance.
These results suggest that the effectiveness of self-training heavily relies on the method used to determine the pseudo-labeling threshold. When the threshold is automatically determined, self-training (i.e. PL ) can perform competitively, indicating that this approach has the potential to improve results compared to the supervised RF.
However, when a fixed threshold is applied, self-training tends to yield inferior results compared to the supervised learning approach. This suggests that an arbitrarily chosen threshold might not effectively capture the underlying patterns in the data for the pseudo-labeling process, leading to suboptimal performance.
Moreover, when the threshold is too low as for $\theta\in\{0.5,0.7\}$, pseudo-labeling is likely to produce label noise and degrade the performance of self-training with respect to the supervised RF classifier in all cases. When the threshold it is too high (i.e. $\theta=0.9$), self-training becomes competitive compared to RF on Isolet and MNIST, but the quantity of pseudo-labeled unlabeled examples seems not to be sufficient to learn efficiently.
In summary, the findings emphasize the importance of a dynamic and adaptive threshold selection mechanism when implementing self-training.